# Diverse Text-to-Image Generation via Contrastive Noise Optimization

### v2

@@ 1 Introduction
In recent years, diffusion models (Ho et al., 2020; Song et al., 2021; Rombach et al., 2022) have emerged as the leading paradigm for text-to-image (T2I) generation. A key driver of their success is the use of text-guided inference, which steers the generation process to produce images that are not only high-fidelity but also closely aligned with a given prompt. To maximize this alignment and enhance image quality, practitioners often employ strong guidance mechanisms, with techniques like Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) becoming a standard practice. However, this pursuit of high fidelity comes at a significant cost: a pronounced lack of diversity. Under strong textual guidance, the model’s outputs often collapse into a few dominant modes, failing to capture the rich variety of interpretations a text prompt can have. This fidelity-diversity trade-off (Dhariwal & Nichol, 2021) remains a critical bottleneck, severely restricting the creative potential of T2I models.
To address this challenge, a common line of work has focused on interventions during the iterative denoising process. These approaches typically optimize intermediate latents (Corso et al., 2024; Kirchhof et al., 2025) or manipulate text embeddings (Sadat et al., 2024; Um & Ye, 2025b) to enforce separation between samples, while other strategies rely on multi-agent systems or complex fine-tuning schedules (Ghosh et al., 2017). While these methods have shown promise, they often provide only modest gains in diversity and introduce significant practical challenges. Their limitations include high implementation complexity, potential instability, and a heavy reliance on fragile hyperparameter tuning, which hinders their robustness and scalability.
@@
In this work, we tackle the problem at its fundamental source by shaping the initial noise distribution. We introduce Contrastive Noise Optimization, a simple yet powerful pre-processing framework that optimizes a batch of initial noise latents to be inherently distinct before the diffusion process begins. Our main contribution is a contrastive objective defined in the data space of denoised predictions via Tweedie’s formula (Robbins, 1992). Specifically, the objective integrates two complementary forces: a repulsion term that pushes Tweedie predictions from different noises apart, maximizing semantic distance to promote diversity; and an attraction term that anchors each optimized noise to its unoptimized counterpart, preserving semantic fidelity and preventing distributional shift.
We highlight that unlike existing approaches that impose diversity loss in the latent space (e.g., the space of $\mathbf{z}_{T}$), our method applies the loss in the denoised Tweedie space, which provides the best estimate of clean generated data, thereby maximizing effectiveness. See Figure 2 for an illustration. To provide theoretical intuitions, we extend the traditional mutual information bound of the InfoNCE loss to incorporate both positive and negative pairs, thereby offering a unified explanation of how our contrastive objective simultaneously promotes diversity and preserves semantic fidelity. Comprehensive experiments confirm that this simple preprocessing substantially enhances the ability of modern T2I models to generate visually distinct images with minimal loss in quality or text–image alignment. We highlight that our method requires neither heavy computation nor extensive hyperparameter tuning, unlike existing approaches (Sadat et al., 2024; Um & Ye, 2025b).
Our key contributions are summarized as follows:
  • We introduce Contrastive Noise Optimization, a simple yet powerful pre-processing framework that optimizes initial noise latents to be inherently distinct in the denoised Tweedie space before the diffusion process begins, tackling the diversity issue at its fundamental source.
  • We provide theoretical insights by extending the mutual information bound of the InfoNCE loss to incorporate both positive and negative pairs, offering a unified explanation of how our contrastive objective simultaneously promotes diversity and preserves semantic fidelity.
  • We show through comprehensive experiments that this lightweight pre-processing significantly boosts visual diversity with minimal quality or text-image alignment loss, while remaining robust to hyperparameters and broadly applicable across modern T2I backbones.
@@ 2 Related Work
Improving the diversity of diffusion models has recently attracted much attention, mainly due to their increasing use in critical applications such as text-to-image generation (Rombach et al., 2022). One prominent effort is CADS (Sadat et al., 2024), which enhances sample diversity by gradually annealing noise perturbations on conditional embeddings. Although effective, their approach is sensitive to the noise annealing schedule and requires laborious hyperparameter searches to see the diversity gain. A fundamentally different strategy is seen in Particle Guidance (PG) (Corso et al., 2024). The idea is to repel intermediate latent samples that share the same condition, thereby encouraging the final generated samples to exhibit distinct features. While it does not require difficult parameter searches like CADS, it often provides limited diversity gain (Kirchhof et al., 2025).s This approach shares a similar spirit as PG (Corso et al., 2024) and incorporates diversity-improving guidance for repelling intermediate latent instances during inference, yet in a sparse manner, i.e., not at every inference timestep. A key distinction from ours is that its diversity optimization (by injecting guidance) is performed over inference time, which may be more expensive compared to ours that focuses on the initial latent space.
@@
A related yet different task is to generate minority samples – low-density instances in the data manifold (Sehwag et al., 2022; Um & Ye, 2023; 2025b). Pioneer works in this area are offered by Sehwag et al. (2022); Um & Ye (2023), which share a similar idea of incorporating classifier guidance (Dhariwal & Nichol, 2021) to push intermediate samples toward low-density regions. The reliance on external classifiers was addressed in Um & Ye (2024; 2025b); Um et al. (2025), offering self-contained approaches for producing minority samples with diffusion models. While relevant, the task of generating low-density minority samples is distinct from improving diversity and does not guarantee distinct outputs. Notably, MinorityPrompt (Um & Ye, 2025b) considers text-to-image generation and provides a prompt optimization framework that can also be used to enhance the diversity of generated samples. However, it requires optimizing the diversity-improving prompt during inference, which imposes substantial computational overhead (Um & Ye, 2025b).
@@
Initial noise optimization in diffusion models has been explored in various contexts (Guo et al., 2024; Ahn et al., 2024). One instance is InitNO (Guo et al., 2024), where the idea is to optimize the initial noise latent to promote improved prompt alignment in text-to-image generation. A distinction with respect to to ours is that their focus is on enhancing text adherence, unlike ours. Another notable work was done by Ahn et al. (2024), who aim to characterize the influence of classifier-free guidance (Ho & Salimans, 2021) through a properly optimized latent noise, enabled by an additional neural network that maps to the optimal noise. While interesting, their focus is inherently distinct from ours. To the best of our knowledge, our framework is the first to incorporate the idea of noise optimization for addressing the diversity challenge of diffusion models.
@@ 3.2 Reverse process and denoising via Tweedie’s formula
The core of this process is the network $\bm{\epsilon}_{\theta}$, which is trained to predict the noise component from the noisy latent vector $\mathbf{z}_{t}$ based on the condition $\mathbf{c}$. The key insight is that this trained network can be used to directly estimate the original clean latent $\mathbf{z}_{0}$ at any timestep $t$. This denoised estimate $\hat{\mathbf{z}}_{0}$ is implemented via Tweedie’s formula (Chung et al., 2025; Um & Ye, 2025a), which for our specific noise model takes the form:
$\hat{\mathbf{z}}_{0}(\mathbf{z}_{t},t,\mathbf{c})\coloneq\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\right).$ (4)
@@ 3.3 Information Noise-Contrastive Estimation (InfoNCE)
Information Noise-Contrastive Estimation (InfoNCE) (van den Oord et al., 2019) is a fundamental objective for self-supervised representation learning (Chen et al., 2020; He et al., 2020). It aims to construct an embedding space that maximizes mutual information (Cover, 1999) between representations of positive (similar) pairs while minimizing it for negative (dissimilar) pairs. The learning process can be viewed as a classification task in which, for a given anchor sample, the model must correctly identify its positive counterpart from a set of negative samples.
Specifically, for an anchor embedding vector $\mathbf{z}_{i}$, its positive pair $\mathbf{z}_{j}$, and a set of B1B-1 negative samples $\{\mathbf{z}_{k}\}_{k=1,k\neq j}^{B}$, the InfoNCE loss is formulated as
$\mathcal{L}_{\text{InfoNCE}}\coloneqq\mathbb{E}_{i}\left[-\log\left(\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1}^{B}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}\right)\right],$ (5)
where $\text{sim}(\cdot,\cdot)$ denotes a similarity measure (e.g., cosine similarity) between two representation vectors, and $\tau$ is a temperature parameter controlling the sharpness of the distribution. Intuitively, the loss encourages the anchor to be close to its positive pair while pushing it away from all negative samples, thereby tightening intra-class similarity and enlarging inter-class separation in the embedding space.
@@ 4.1 Optimizing initial noise with contrastive loss between Tweedies
While text-to-image diffusion models generate high-quality images for a given prompt $c$, the standard approach of sampling a random initial noise $\mathbf{z}_{T}$ often yields outputs that collapse into similar modes. This tendency is a primary obstacle to achieving a diverse set of generations. To address this issue, we propose a novel approach that enhances diversity by optimizing the initial noise $\mathbf{z}_{T}$ itself, prior to commencing the standard DDIM sampling process. This procedure is detailed in Algorithm 1 in Appendix B.1.
The algorithm proceeds as follows. First, we sample a batch of initial latent codes $\mathcal{B}=\{\mathbf{z}^{i}_{T}\}_{i=1}^{B}$ from a standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$. Immediately after, we compute the initial target latents, $\{\mathbf{z}^{i}_{0|T}\}^{B}_{i=1}$, by applying the denoising estimator defined in Equation Eq. (4) to this initial noise at timestep $T$. Each resulting $\mathbf{z}^{i}_{0|T}$ is therefore the model’s one-step prediction of the clean latent $\mathbf{z}_{0}$ from the noise $\mathbf{z}^{i}_{T}$. These pre-computed latents then serve as fixed anchors, each defining a unique identity for its respective sample throughout the optimization.
Before computing the loss, we employ a practical optimization to enhance efficiency. The high dimensionality of the latents $(B,C,S,S)$ makes the pairwise similarity calculation computationally intensive. We found experimentally that applying an adaptive average pooling operation to downsample the latents to a sufficiently smaller spatial resolution $(B,C,w,w)$, where $w<S$, did not compromise performance (Section 5.1). This step substantially reduces memory usage and accelerates the similarity matrix computation, making the optimization process more tractable. Downsampled latents $\{\mathbf{z}^{i,\text{fixed}}_{0|T}\}^{B}_{i=1}$, $\{\mathbf{z}^{i,\text{opt}}_{0|T}\}^{B}_{i=1}$ are then normalized, and the noise $\{\mathbf{z}^{i}_{T}\}_{i=1}^{B}$ is updated using a contrastive loss $\mathcal{L_{\text{CNO}}}$:
$\mathcal{L}_{\text{CNO}}\coloneqq\mathbb{E}_{i\sim\mathcal{B}}\left[-\log\frac{\exp(\text{sim}(\mathbf{z}_{0|T}^{i,\text{opt}},\mathbf{z}_{0|T}^{i,\text{fixed}})/\tau)}{\sum_{j\in\mathcal{B}}\exp(\text{sim}(\mathbf{z}_{0|T}^{i,\text{opt}},\mathbf{z}_{0|T}^{j,\text{opt}})/\tau)}\right].$ (6)
@@
Attraction (Numerator). It encourages the current latent $\mathbf{z}^{i,\text{opt}}_{0|T}$ to remain similar to its corresponding initial target latent $\mathbf{z}^{i,\text{fixed}}_{0|T}$. This ensures that each sample maintains coherence with its initial concept and does not drift away during optimization.
Repulsion (Denominator). It pushes the current latent $\mathbf{z}^{i,\text{opt}}_{0|T}$ to be dissimilar from all other current latents $\{\mathbf{z}_{0|T}^{j,\text{opt}}\}^{B}_{j=1,j\neq i}$ in the batch. This directly promotes diversity by forcing the latent representations to disperse within the batch.
@@
Stop-gradient for computational efficiency. Optimizing the loss in Eq. (6) requires backpropagation through diffusion models, which can incur substantial computational overhead. To mitigate this, we apply a stopgrad operator (Chen & He, 2021) on the model path used in computing the Tweedie’s estimate. As also demonstrated in Ahn et al. (2024), this simple strategy yields significant savings in training cost with only marginal impact on performance (see Table 4).

### v3

@@ 1 Introduction
In recent years, diffusion models (Ho et al., 2020; Song et al., 2021; Rombach et al., 2022) have emerged as the leading paradigm for text-to-image (T2I) generation. A key driver of their success is the use of text-guided inference, which steers the generation process to produce images that are not only high-fidelity but also closely aligned with a given prompt. To maximize this alignment and enhance image quality, practitioners often employ strong guidance mechanisms, with techniques like Classifier-Free Guidance (CFG) (Ho and Salimans, 2021) becoming a standard practice. However, this pursuit of high fidelity comes at a significant cost: a pronounced lack of diversity. Under strong textual guidance, the model’s outputs often collapse into a few dominant modes, failing to capture the rich variety of interpretations a text prompt can have. This fidelity-diversity trade-off (Dhariwal and Nichol, 2021) remains a critical bottleneck, severely restricting the creative potential of T2I models.
To address this challenge, a common line of work has focused on interventions during the iterative denoising process. These approaches typically optimize intermediate latents (Corso et al., 2024; Kirchhof et al., 2025) or manipulate text embeddings (Sadat et al., 2024; Um and Ye, 2025b) to enforce separation between samples, while other strategies rely on multi-agent systems or complex fine-tuning schedules (Ghosh et al., 2017). Although these methods have shown promise, they often require repeated adjustments during sampling and thus suffer from notable limitations, including substantial computational overhead (e.g., DiversityPrompt (Um and Ye, 2025b); see Table 8) and limited gains under accelerated few-step models, where opportunities for intervention are intrinsically scarce.
@@
In this work, we tackle the diversity problem at its fundamental source by shifting the paradigm from inference-time interventions to initial-noise selection. We introduce Contrastive Noise Optimization (CNO), a simple yet powerful one-shot preprocessing framework that optimizes a batch of initial noise latents before sampling begins. Unlike prior methods that intervene repeatedly during the denoising trajectory, CNO performs a single, lightweight optimization step on initial noise and requires no modifications or adjustments during sampling. The key idea is to incorporate a contrastive objective to shape a diversity-encouraging initial noise distribution, drawing inspiration from the structure of InfoNCE-based contrastive learning (van den Oord et al., 2019). Building on this foundation, we introduce structural modifications tailored to the unique demands of diverse T2I generation, enabling diversity to be enhanced while retaining strong semantic fidelity.
At the core of our framework lies a balance between two complementary forces: an attraction term that anchors each optimized noise to its original counterpart, and a repulsion term that encourages semantic separation across samples. Crucially, we impose this contrastive structure not in the raw noise latent space but in the Tweedie denoised prediction space, which provides the diffusion model’s best estimate of clean data. Operating in this space allows the optimization to act directly on meaningful semantic signals, making the refinement of initial noise substantially more effective.
To further control this interplay, we introduce a novel balancing parameter $\gamma$, which modulates the relative strength of attraction and repulsion. Its behavior is analytically grounded through an extended mutual-information perspective (Proposition 2), revealing how $\gamma$ governs the contributions of positive and negative pairs. We additionally provide a closed-form selection rule for $\gamma$ that ensures robust behavior across batch sizes. Practical components such as adaptive latent pooling and a stop-gradient mechanism further enhance the efficiency and scalability of CNO for modern T2I backbones.
Comprehensive experiments demonstrate that our lightweight preprocessing substantially improves diversity across modern T2I frameworks (including Stable Diffusion (Rombach et al., 2022)) while maintaining high image quality and text alignment. Notably, CNO delivers consistent improvements even under accelerated few-step samplers such as FLUX (Labs et al., 2025) and SDXL-Lightning (Lin et al., 2024), where existing diversity approaches exhibit limited gains (see Section D.1).
  • We introduce a paradigm shift from inference-time interventions to initial-noise selection, addressing the diversity problem at its source while eliminating repeated per-step adjustments and remaining effective even in accelerated few-step samplers.
  • We develop a contrastive noise optimization framework tailored for diverse T2I generation, featuring the balancing parameter $\gamma$, an extended mutual-information analysis, and efficient heuristics such as adaptive latent pooling and stop-gradient.
  • We demonstrate state-of-the-art diversity–quality trade-offs across major T2I backbones, including SD1.5, SDXL, SD3, and fast few-step samplers such as FLUX and SDXL-Lightning.
@@ 2 Related Work
Improving the diversity of diffusion models has recently attracted much attention, mainly due to their increasing use in critical applications such as text-to-image generation (Rombach et al., 2022). One prominent effort is CADS (Sadat et al., 2024), which enhances sample diversity by gradually annealing noise perturbations on conditional embeddings. Although effective, their approach is sensitive to the noise annealing schedule and requires laborious hyperparameter searches to see the diversity gain. A fundamentally different strategy is seen in Particle Guidance (PG) (Corso et al., 2024). The idea is to repel intermediate latent samples that share the same condition, thereby encouraging the final generated samples to exhibit distinct features. While it does not require difficult parameter searches like CADS, it often provides limited diversity gains (Kirchhof et al., 2025). This approach shares a similar spirit as PG (Corso et al., 2024) and incorporates diversity-improving guidance for repelling intermediate latent instances during inference, yet in a sparse manner, i.e., not at every inference timestep. A key distinction from ours is that its diversity optimization (by injecting guidance) is performed over inference time, which may be more expensive compared to ours that focuses on the initial latent space.
@@
A related yet different task is to generate minority samples – low-density instances in the data manifold (Sehwag et al., 2022; Um and Ye, 2023; 2025b). Pioneer works in this area are offered by Sehwag et al. (2022); Um and Ye (2023), which share a similar idea of incorporating classifier guidance (Dhariwal and Nichol, 2021) to push intermediate samples toward low-density regions. The reliance on external classifiers was addressed in Um and Ye (2024; 2025b); Um et al. (2025), offering self-contained approaches for producing minority samples with diffusion models. While relevant, the task of generating low-density minority samples is distinct from improving diversity and does not guarantee distinct outputs. Notably, MinorityPrompt (Um and Ye, 2025b) considers text-to-image generation and provides a prompt optimization framework that can also be used to enhance the diversity of generated samples. However, it requires optimizing the diversity-improving prompt during inference, which imposes substantial computational overhead (Um and Ye, 2025b).
@@
Initial noise optimization in diffusion models has been explored in various contexts (Guo et al., 2024; Ahn et al., 2024). One instance is InitNO (Guo et al., 2024), where the idea is to optimize the initial noise latent to promote improved prompt alignment in text-to-image generation. A distinction with respect to ours is that their focus is on enhancing text adherence, unlike ours. Another notable work was done by Ahn et al. (2024), who aim to characterize the influence of classifier-free guidance (Ho and Salimans, 2021) through a properly optimized latent noise, enabled by an additional neural network that maps to the optimal noise. While interesting, their focus is inherently distinct from ours. To the best of our knowledge, our framework is the first to incorporate the idea of noise optimization for addressing the diversity challenge of diffusion models.
@@ 3.2 Reverse Process and Denoising via Tweedie’s Formula
The core of this process is the network $\bm{\epsilon}_{\theta}$, which is trained to predict the noise component from the noisy latent vector $\mathbf{z}_{t}$ based on the condition $\mathbf{c}$. The key insight is that this trained network can be used to directly estimate the original clean latent $\mathbf{z}_{0}$ at any timestep $t$. This denoised estimate $\hat{\mathbf{z}}_{0}$ is implemented via Tweedie’s formula (Chung et al., 2025; Um and Ye, 2025a), which for our specific noise model takes the form:
${\hat{\mathbf{z}}_{0|t}}(\mathbf{z}_{t},t,\mathbf{c})\coloneq\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\right).$ (4)
@@ 3.3 Information Noise-Contrastive Estimation (InfoNCE)
Information Noise-Contrastive Estimation (InfoNCE) (van den Oord et al., 2019) is a fundamental objective for self-supervised representation learning (Chen et al., 2020; Kim et al., 2021; Jang et al., 2023; Kim et al., 2025). It aims to construct an embedding space that maximizes mutual information (Cover, 1999) between representations of positive (similar) pairs while minimizing it for negative (dissimilar) pairs. The learning process can be viewed as a classification task in which, for a given anchor sample, the model must correctly identify its positive counterpart from a set of negative samples.
Specifically, for an anchor embedding vector $\mathbf{z}_{i}$, its positive pair $\mathbf{z}^{\text{pos}}_{i}$, and a set of B1B-1 negative samples $\{\mathbf{z}_{j}\}_{j=1,j\neq i}^{B}$, the InfoNCE loss is formulated as
$\mathcal{L}_{\text{InfoNCE}}\coloneqq\frac{1}{B}\sum_{i=1}^{B}\left[-\log\left(\frac{\exp(f(\mathbf{z}_{i},\mathbf{z}^{\text{pos}}_{i})/\tau)}{\sum_{j=1}^{B}\exp(f(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}\right)\right],$ (5)
where $f(\cdot,\cdot)$ denotes a similarity measure (e.g., cosine similarity) between two representation vectors, and $\tau$ is a temperature parameter controlling the sharpness of the distribution. Intuitively, the loss encourages the anchor to be close to its positive pair while pushing it away from all negative samples, thereby tightening intra-class similarity and enlarging inter-class separation in the embedding space.
@@ 4.1 Optimizing Initial Noise with Contrastive Loss between Tweedies
A core challenge in text-to-image diffusion models is that independently sampled initial noises $\mathbf{z}_{T}$ often lead to generations that collapse into similar modes, even under varied stochasticity. Rather than intervening during the denoising trajectory as in prior diversity methods, we enhance diversity at its fundamental source by optimizing the initial noises themselves before sampling begins. Our approach, which we call Contrastive Noise Optimization (CNO), refines a batch of initial noises using a contrastive objective applied in the Tweedie denoised prediction space, enabling the noises to be semantically well-separated while remaining faithful to their original distribution. The full procedure is provided in Algorithm 1 in Section B.1.
The algorithm proceeds as follows. First, we sample a batch of initial latent codes ${\mathbf{Z}_{T}}=\{\mathbf{z}^{i}_{T}\}_{i=1}^{B}$ from a standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$. Hereafter, we denote the denoised estimate from Equation Eq. (4) as $\hat{\mathbf{z}}_{0|t}$. Using this, we compute the initial target latents, $\{\hat{\mathbf{z}}^{i}_{0|T}\}^{B}_{i=1}$, by applying the denoising estimator defined in Equation Eq. (4) to this initial noise at timestep $T$. Each resulting $\hat{\mathbf{z}}^{i}_{0|T}$ is therefore the model’s one-step prediction of the clean latent $\mathbf{z}_{0}$ from the noise $\mathbf{z}^{i}_{T}$. These pre-computed latents then serve as fixed anchors, each defining a unique identity for its respective sample throughout the optimization.
Before computing the loss, we employ a practical optimization to enhance efficiency. The high dimensionality of the latents $(B,C,S,S)$ makes the pairwise similarity calculation computationally intensive. We found experimentally that applying an adaptive average pooling operation to downsample the latents to a sufficiently smaller spatial resolution $(B,C,w,w)$, where $w<S$, did not compromise performance (Section 5.1). This step substantially reduces memory usage and accelerates the similarity matrix computation, making the optimization process more tractable. Downsampled latents $\{\hat{\mathbf{z}}^{i,\text{ref}}_{0|T}\}^{B}_{i=1}$, $\{\hat{\mathbf{z}}^{i}_{0|T}\}^{B}_{i=1}$ are then normalized, and the noise $\{\mathbf{z}^{i}_{T}\}_{i=1}^{B}$ is updated using a contrastive loss $\mathcal{L_{\text{CNO}}}$:
$\mathcal{L}_{\text{CNO}}\coloneqq\frac{1}{B}\sum_{i=1}^{B}\left[-\log\left(\frac{\exp(f(\hat{\mathbf{z}}_{0|T}^{i},\hat{\mathbf{z}}_{0|T}^{i,\text{ref}})/\tau)}{\sum_{j=1}^{B}\exp(f(\hat{\mathbf{z}}_{0|T}^{i},\hat{\mathbf{z}}_{0|T}^{j})/\tau)}\right)\right].$ (6)
@@
Attraction (Numerator). It encourages the current latent $\hat{\mathbf{z}}^{i}_{0|T}$ to remain similar to its corresponding initial target latent $\hat{\mathbf{z}}^{i,\text{ref}}_{0|T}$. This ensures that each sample maintains coherence with its initial concept and does not drift away during optimization.
Repulsion (Denominator). It pushes the current latent $\hat{\mathbf{z}}^{i}_{0|T}$ to be dissimilar from all other current latents $\{\hat{\mathbf{z}}_{0|T}^{j}\}^{B}_{j=1,j\neq i}$ in the batch. This directly promotes diversity by forcing the latent representations to disperse within the batch.
@@
Stop-gradient for computational efficiency. Optimizing the loss in Eq. (6) requires backpropagation through diffusion models, which can incur substantial computational overhead. To mitigate this, we apply a stopgrad operator (Chen and He, 2021) on the model path used in computing the Tweedie’s estimate. As also demonstrated in Ahn et al. (2024), this simple strategy yields significant savings in training cost with only marginal impact on performance (see Table 8).